# Algorithmic Collusion at Test Time: A Meta-game Design and Evaluation

### v1

@@ Abstract
The threat of algorithmic collusion, and whether it merits regulatory intervention, remains debated, as existing evaluations of its emergence often rely on long learning horizons, assumptions about counterparty rationality in adopting collusive strategies, and symmetry in hyperparameters and economic settings among players. To study collusion risk, we introduce a meta-game design for analyzing algorithmic behavior under test-time constraints. We model agents as possessing pretrained policies with distinct strategic characteristics (e.g., competitive, naively cooperative, robustly collusive), and formulate the problem as selecting a meta-strategy that combines a pretrained, initial policy with an in-game adaptation rule. We seek to examine whether collusion can emerge under rational choices and how agents co-adapt toward cooperation or competition. To this end, we sample normal-form empirical games over meta-strategy profiles, compute relevant game statistics (e.g., payoffs against individuals and regret against an equilibrium mixture of opponents), and construct empirical best-response graphs to uncover strategic relationships. We evaluate both reinforcement-learning and LLM-based strategies in repeated pricing games under symmetric and asymmetric cost settings, and present findings on the feasibility of algorithmic collusion and the effectiveness of pricing strategies in practical “test-time” environments. The source code and the full paper with appendix are available at: https://github.com/chailab-rutgers/CollusionMetagame.
@@ 1 Introduction
In this work, we seek to evaluate the viability and effectiveness of algorithmic collusion in “test-time” settings, where an agent must adapt within a limited number of interactions to opponents whose policies and cost structures may differ from those encountered during training. To this end, we propose a meta-game framework. Different learning algorithms (e.g., Q-learning, no-regret methods, and LLMs) are first used at training time to generate candidate initial policies for deployment. These policies are then grouped into strategic categories based on performance metrics, including their propensity to collude with a training-time partner and their robustness against a best-response opponent. At test time, a strategy is formed by pairing an initial policy sampled from a selected category with an adaptation rule (e.g., a learning rate) drawn from a set of feasible update procedures for repeated play. A meta-strategy governs both the choice of initial policy category and the manner in which the policy adapts during interaction.
@@ 1.1 Related Work
Collusive behavior has been documented in e-commerce and dynamic pricing settings [ezrachi_algorithmic_2015; Byrne2019Learning; musolff_algorithmic_2022; Assad2024Algorithmic; gianluca2022]. Following Calvano2020Artificial, a growing body of literature investigates algorithmic collusion in laboratory across settings ranging from stylized games, such as the repeated Prisoner’s Dilemma [Meylahn2022Limiting; Wolfram2023Intrinsic], to more realistic auctions [Banchio2022Artificial]. These studies also consider different algorithms, including reinforcement learning [klein_autonomous_2021; hettich_algorithmic_2021], bounded-regret methods [Hansen2021Frontiers], and LLMs [Fish2024Algorithmic; agrawal2025evaluatingllmagentcollusion]. Several follow-up studies suggest that prior findings may overstate the threat of algorithmic collusion, as they rely on restrictive modeling assumptions [abada_algorithmic_2024]. Collusion typically emerges after long training horizons—on the order of millions of repeated-game rounds—and tends to depend on symmetric algorithmic configurations, suggesting implicit coordination or prior communication [lambin_less_2023]. Moreover, the algorithmic configurations that give rise to collusion may be irrational in practice, being outperformed by more robust or commonly used alternatives [Abada2023Artificial; bichler2024onlineoptimizationalgorithmsrepeated]. Thus, the emergence of algorithmic tacit collusion among rational agents under realistic conditions and finite interaction horizons remains an open question.
@@ 2 Preliminaries
If both policies are deterministic at state $(p_{j},p_{-j})$, e.g., $\pi_{j}((p_{j},p_{-j}),p_{j})=1$ and $\pi_{-j}((p_{j},p_{-j}),p_{-j})=1$, then $V^{\pi_{j}|\pi_{-j}}((p_{j},p_{-j}))=\frac{r_{j}((p_{j},p_{-j}))}{1-\gamma}$.
@@ 3.1.3 Meta-strategy
Intuitively, the relative magnitudes of $\overline{V}^{\pi_{j}|\pi_{b}}$ and $\overline{V}^{\pi_{b}|\pi_{j}}$ characterize the strategic nature of the interaction between the two policies. When these values are highly asymmetric, it indicates an exploitative relationship, typically with $\pi_{b}$ exploiting $\pi_{j}$, which reflects $\pi_{j}$’s lack of robustness. When both values are low, the interaction tends to settle in a competitive absorbing state, indicating robustness but limited cooperative gains. In contrast, when both values are high and comparable, the two policies exhibit robust cooperation, achieving mutually beneficial outcomes while remaining stable against unilateral deviations. [^2] See Fig. 5 in Appendix for examples of Q-learning policies in different strategic categories.
@@
Figure 1: A toy meta-game for repeated Prisoner’s Dilemma on canonical strategies. The NE of this meta-game leads to cooperation. [^4] The CR and PC are evaluated with the discount factor $\gamma=0.95$. Q-learning is adopted for both pretraining and test time. We adopt the normalization of the initial Q-values following the procedure described in Appendix A.2 and set exploration $\varepsilon=0$. We stop at $t=80$ because all base games are converged.See Example 3.1 for more discussion.
@@
We categorize these policies. All policies achieve high PC with their respective pretrained partners due to cooperative behavior. However, WSLS and AC both yield low mean state values against their best responses, resulting in highly asymmetric CR, as they can be exploited by a best-response policy (e.g., always Defect). By contrast, TfT attains high CR, as its best response is to always cooperate.
@@ 4.1.1 Meta-strategies
Tertiles are computed within each pretraining algorithm. We also consider a baseline category of randomly initialized (RD) policies. Each initial policy category is paired with a set of algorithm-specific update rules spanning a range of adaptation speeds, from fast to slow. We discuss Q-learning in Sec. 4.2, UCB in Sec. 4.3 and LLM in Sec. 4.4. For all meta-game evaluations, additional results including the payoff matrices the best-response graphs are referenced in Appendix A.6 for Q-learning, Appendix B.2 for UCB and Appendix C.3 for LLM.
@@ 4.2 Tabular Q-learning for Pricing
$\phi(Z_{t})=\begin{cases}\arg\max_{p_{j}\in\mathcal{P}}Z_{t}(s,p_{j})&\text{ w.p. }1-\varepsilon\\ p_{j}\sim\text{Unif}(\mathcal{P}_{j})&\text{ otherwise.}\end{cases}\;\;\text{ for each }s\in\mathcal{S}.$ (2)
@@ 4.2.1 Symmetric Costs
Table 10 and Fig. 3 present our main findings for symmetric base games, evaluated at t=10,000t=10,000.111111Based on Amazon’s 15-minute pricing updates, $t=3,000$ is approximately one month, and 10,000 approximate three months. https://sell.amazon.com/tools/automate-pricing We identify a PSNE at (C, 0.5) and a max-entropy MSNE involving three meta-strategies: (C, 0.5), (RC, 0.05), and (RC, 0.005).
@@ 4.2.2 Asymmetric Costs
We construct a meta-game where the set of meta-strategies includes a mixture of rescaling factors, $f\in\{0.5,1.0\}$.121212f=0f=0 is excluded due to the very limited number of updates it induces. Due to the large strategy space, we focus on competitive meta-strategies that appear in the equilibria of their respective single-rescaling-factor meta-games. The base game uses symmetric costs of one, evaluated at $t=10,000$.
@@ 4.3 UCB for Pricing
$\phi(Z_{t})$ $=\arg\max_{p\in\mathcal{P}}\text{UCB}_{j,t}(s,p)\text{ for all $s\in\mathcal{S}$ where }$ (4)
$\text{UCB}_{j,t}(s,p)$ $=\frac{1}{\text{count}_{t}(s,p)}r_{t}(s,p)+\sqrt{\frac{2\ln\text{count}_{t}(s)}{\text{count}_{t}(s,p)}}$ (5)
$\text{ and }\text{count}_{t}(s)$ $=\sum_{p\in\mathcal{P}}\text{count}_{t}(s,p)$ (6)
@@ 4.4.1 Meta-game Evaluation
We focus on strategies generated by GPT5-mini, since GPT5-nano mostly converges immediately to the competitive price regardless of the paired opponent. Specifically, we select six meta-strategies that roughly span LC (p0h0, p1h2 and p2h1), C (p2h3, p1h1), and RC (p0h2). Table 4 and Appendix Fig. 25 summarize main findings for the symmetric repeated pricing game with $c_{1}=c_{2}=1$, evaluated at $t=50$ with $N_{\text{discrete}}=4,N_{\text{base}}=40$. [^13] We observe that all policies stop updating after about 30 rounds. Prices are discretized into four levels due to API costs, and initial states are uniformly sampled for the base game.
@@ 5 Discussion
Natural extensions of our framework include modeling heterogeneous beliefs on the opponent’s cost or representation initialization and analyze through Bayesian Nash equilibrium, as well as broadening meta-strategy coverage, particularly for LLMs with richer prompt designs and pretraining histories. Additionally, cross-algorithm meta-games could shed further light on how algorithmic diversity affects the emergence and stability of collusion.
@@ A.2 Normalization of Q-values
$+\bigg(\tilde{Z}_{j}(s,p)-\min_{p^{\prime}}\tilde{Z}_{j}(s,p^{\prime})\bigg)\times\frac{V^{\pi_{j}|\pi^{\prime}_{j}}(s)-\min_{p^{\prime}}Q^{\pi_{j}|\pi^{\prime}_{j}}(s,p^{\prime})}{\max_{p^{\prime}}\tilde{Z}_{j}(s,p^{\prime})-\min_{a^{\prime}}\tilde{Z}_{j}(s,p^{\prime})}$ (10)
@@
After normalization, the original argmax of $\tilde{Z}_{j}$ is preserved: $\tilde{Z}_{j}(s,\pi_{j}(s))=V^{\pi_{j}|\pi^{\prime}_{j}}(s)+Q^{N}+f\cdot(Q^{M}-Q^{N})-\max_{s^{*}}V^{\pi_{j}|\pi^{\prime}_{j}}(s^{*})$, and the minimum satisfies $\min_{p^{\prime}}\tilde{Z}_{j}(s,p)=\min_{p^{\prime}}Q^{\pi_{j}|\pi^{\prime}_{j}}(s,p^{\prime})+Q^{N}+f\cdot(Q^{M}-Q^{N})-\max_{s^{*}}V^{\pi_{j}|\pi^{\prime}_{j}}(s^{*})$. The maximum Q-value (at the argmax action) maps to the true value function $V^{\pi_{j}|\pi^{\prime}_{j}}(s)$ and the minimum Q-value maps to $\min_{p^{\prime}}Q^{\pi_{j}|\pi^{\prime}_{j}}(s,p^{\prime})$.
@@ A.3 Categorizing Q-learning Pretrained Policies via PC and CR
Figure 6: PC (Def. 3.1) and CR (Def. 3.2) of 1000 pretrained Q-learning policies with 4 and 15 discretized actions, respectively, under the experiment setting described in Sec. 4.2.
@@ A.6 Additional Results on Q-learning-based Meta-strategies
(b) Change in PC over time. Overall, PC improves over the course of the test time. For RD strategies, although their own $V$ improves, their opponent’s does not, as the value against RD is already high at initialization. For both LC and RC, larger learning rates tend to benefit their opponents more, even when their own $\Delta(V^{\pi_{j}|\pi_{-j}})$ is comparable to that under smaller learning rates. This suggests that LC and RC with large learning rates provide opportunities for cooperation.
@@
Figure 10: The best response graph for Q-learning.
Figure 11: $\Delta_{t}$(CR) and $\Delta_{t}$(PC) for Q-learning.
@@
Figure 12: The best-response graph for Q-learning under symmetric costs $c_{1}=c_{2}=1$ with optimistic initialization $f=1$ and a time horizon of $t=3,000$.
@@
Table 11: The average payoffs per strategy pair of Q-learning under asymmetric costs, $c_{1}=1$ and $c_{2}=0.8$, and $t=10,000,f=1$. Standard errors are omitted as they are negligible in the order of $10^{-4}$. For reference, in this setting, the competitive and monopoly payoffs for player 1 are $\bar{r}_{c=1.0}^{N}=0.17$ and $\bar{r}_{c=1.0}^{M}=0.30$, and for player 2 are $\bar{r}_{c=1.0}^{N}=0.31$ and $\bar{r}_{c=0.8}^{M}=0.44$.
@@
Figure 14: The best-response graph. The directed edge $(u\rightarrow v)$ corresponds to $c_{u}=1.0$ and $c_{v}=0.8$.
Figure 15: The best-response graph. The directed edge $(u\rightarrow v)$ corresponds to $c_{u}=0.8$ and $c_{v}=1.0$.
Figure 16: $\Delta_{t}$(CR) and $\Delta_{t}$(PC) for Q-learning. The high cost strategies are plotted with fillings.
Symmetric costs $c_{1}=c_{2}=1$ with either optimistic or pessimistic initializations ($f=1$ and $f=0.5$) under time horizon $t=10,000$.
@@
Table 12: The average payoffs per strategy pair of Q-learning under symmetric costs $c_{1}=c_{2}=1$ with either optimistic initialization $f=1$ or pessimistic initialization $f=0.5$. Time horizon $t=10,000$. For reference, in this setting, the competitive and monopoly payoffs are $\bar{r}^{N}=0.22$ and $\bar{r}^{M}=0.34$.
@@
Figure 17: The best response graph for Q-learning.
[…]

### v2

@@ Abstract
The threat of algorithmic collusion, and whether it merits regulatory intervention, remains debated, as existing evaluations of its emergence often rely on long learning horizons, assumptions about counterparty rationality in adopting collusive strategies, and symmetry in hyperparameters and economic settings among players. To study collusion risk, we introduce a meta-game design for analyzing algorithmic behavior under test-time constraints. We model agents as possessing pretrained policies with distinct strategic characteristics (e.g., competitive, naively cooperative, or robustly collusive), and formulate the problem as selecting a meta-strategy that combines a pretrained, initial policy with an in-game adaptation rule. We seek to examine whether collusion can emerge under rational choices and how agents co-adapt toward cooperation or competition. To this end, we sample normal-form empirical games over meta-strategy profiles, compute relevant game statistics (e.g., payoffs against individuals and regret against an equilibrium mixture of opponents), and construct empirical best-response graphs to uncover strategic relationships. We evaluate reinforcement-learning, UCB, and LLM-based strategies in repeated pricing games under symmetric and asymmetric cost settings, and present findings on the feasibility of algorithmic collusion and the effectiveness of pricing strategies in practical “test-time” environments. The source code is available at: https://github.com/chailab-rutgers/CollusionMetagame.
@@ 1 Introduction
In this work, we seek to evaluate the viability and effectiveness of algorithmic collusion in “test-time” settings, where an agent must adapt within a limited number of interactions to opponents whose policies and economic settings (e.g., cost or quality) may differ from those encountered during training. To this end, we propose a meta-game framework. Different learning algorithms (e.g., Q-learning, no-regret methods, and LLMs) are first used at training time to generate candidate initial policies for deployment. These policies are then grouped into strategic categories based on performance metrics, including their propensity to collude with a training-time partner and their robustness against a best-response opponent. At test time, a strategy is formed by pairing an initial policy sampled from a selected category with an adaptation rule (e.g., a learning rate) drawn from a set of feasible update procedures for repeated play. A meta-strategy governs both the choice of initial policy category and the manner in which the policy adapts during interaction.
@@ 1.1 Related Work
Collusive behavior has been documented in e-commerce and dynamic pricing settings [ezrachi_algorithmic_2015; Byrne2019Learning; musolff_algorithmic_2022; Assad2024Algorithmic; gianluca2022]. Following Calvano2020Artificial, a growing body of literature investigates algorithmic collusion in the laboratory across settings ranging from stylized games, such as the repeated Prisoner’s Dilemma [Meylahn2022Limiting; Wolfram2023Intrinsic], to more realistic auctions [Banchio2022Artificial]. These studies also consider different algorithms, including reinforcement learning [klein_autonomous_2021; hettich_algorithmic_2021], bounded-regret methods [Hansen2021Frontiers], and LLMs [Fish2024Algorithmic; agrawal2025evaluatingllmagentcollusion]. Several follow-up studies suggest that prior findings may overstate the threat of algorithmic collusion, as they rely on restrictive modeling assumptions [abada_algorithmic_2024]. Collusion typically emerges after long training horizons—on the order of millions of repeated-game rounds—and tends to depend on symmetric algorithmic configurations, suggesting implicit coordination or prior communication [lambin_less_2023]. Moreover, the algorithmic configurations that give rise to collusion may be irrational in practice, being outperformed by more robust or commonly used alternatives [Abada2023Artificial; bichler2024onlineoptimizationalgorithmsrepeated]. Thus, the emergence of algorithmic tacit collusion among rational agents under realistic conditions and finite interaction horizons remains an open question.
@@ 2 Preliminaries
If state $(p_{j},p_{-j})$ is an absorbing state, i.e., $\pi_{j}((p_{j},p_{-j}),p_{j})=1$ and $\pi_{-j}((p_{j},p_{-j}),p_{-j})=1$, then $V^{\pi_{j}|\pi_{-j}}((p_{j},p_{-j}))=\frac{r_{j}((p_{j},p_{-j}))}{1-\gamma}$.
@@ 3.1.3 Meta-strategy
Intuitively, the relative magnitudes of $\overline{V}^{\pi_{j}|\pi_{b}}$ and $\overline{V}^{\pi_{b}|\pi_{j}}$ characterize the strategic nature of the interaction between the two policies. When these values are highly asymmetric, it indicates an exploitative relationship, typically with $\pi_{b}$ exploiting $\pi_{j}$, which reflects $\pi_{j}$’s lack of robustness. When both values are low, the interaction tends to settle in a competitive absorbing state, indicating robustness but limited cooperative gains. In contrast, when both values are high and comparable, the two policies exhibit robust cooperation, achieving mutually beneficial outcomes while remaining stable against unilateral deviations. [^2] See Fig. 5 in the appendix for examples of Q-learning policies in different strategic categories.
@@
Figure 1: A toy meta-game for repeated Prisoner’s Dilemma on canonical strategies. The NE of this meta-game leads to cooperation. See Example 3.1 for more discussion.
@@
We categorize these policies. All policies achieve high PC with their respective pretrained partners due to cooperative behavior. However, WSLS and AC both yield low mean state values against their best responses, resulting in highly asymmetric CR, as they can be exploited by a best-response policy (e.g., Always-Defect). By contrast, TfT attains high CR, as its best response is to always cooperate.
@@ 4.1.1 Meta-strategies
Tertiles are computed within each pretraining algorithm. We also consider a baseline category of randomly initialized (RD) policies. Each initial policy category is paired with a set of algorithm-specific update rules spanning a range of adaptation speeds, from fast to slow. We discuss Q-learning in Sec. 4.2, UCB in Sec. 4.3 and LLM in Sec. 4.4. For all meta-game evaluations, additional results including the payoff matrices and the best-response graphs are referenced in Appendix A.6 for Q-learning, Appendix B.2 for UCB, and Appendix C.3 for LLM.
@@ 4.2 Tabular Q-learning for Pricing
$\phi(Z_{t})=\begin{cases}\arg\max_{p_{j}\in\mathcal{P}_{j}}Z_{t}(s,p_{j})&\text{ w.p. }1-\varepsilon\\ p_{j}\sim\text{Unif}(\mathcal{P}_{j})&\text{ otherwise.}\end{cases}\;\;\text{ for each }s\in\mathcal{S}.$ (2)
@@ 4.2.1 Symmetric Costs
Table 8 and Fig. 3 present our main findings for symmetric base games, evaluated at t=10,000t=10,000.999Based on Amazon’s 15-minute pricing updates, $t=3,000$ is approximately one month, and $t=10,000$ approximates three months. https://sell.amazon.com/tools/automate-pricing We identify a PSNE at (C, 0.5) and a max-entropy MSNE involving three meta-strategies: (C, 0.5), (RC, 0.05), and (RC, 0.005).
@@ 4.2.2 Asymmetric Costs
We construct a meta-game where the set of meta-strategies includes a mixture of rescaling factors, $f\in\{0.5,1.0\}$.101010f=0f=0 is excluded due to the very limited number of updates it induces. Due to the large strategy space, we focus on competitive meta-strategies that appear in the equilibria of their respective single-rescaling-factor meta-games. The base game uses symmetric costs of one, evaluated at $t=10,000$.
@@ 4.3 UCB for Pricing
$\phi(Z_{t})$ $=\arg\max_{p_{j}\in\mathcal{P}_{j}}\text{UCB}_{j,t}(s,p_{j})\text{ for all $s\in\mathcal{S}$ where }$ (4)
$\text{UCB}_{j,t}(s,p_{j})$ $=\frac{1}{\text{count}_{t}(s,p_{j})}r_{t}(s,p_{j})+\sqrt{\frac{2\ln\text{count}_{t}(s)}{\text{count}_{t}(s,p_{j})}}$ (5)
$\text{ and }\text{count}_{t}(s)$ $=\sum_{p_{j}\in\mathcal{P}_{j}}\text{count}_{t}(s,p_{j})$ (6)
@@ 4.4.1 Meta-game Evaluation
We focus on strategies generated by GPT5-mini, since GPT5-nano mostly converges immediately to the competitive price regardless of the paired opponent. Specifically, we select six meta-strategies that roughly span LC (p0h0, p1h2 and p2h1), C (p2h3, p1h1), and RC (p0h2). Table 4 and Appendix Fig. 25 summarize main findings for the symmetric repeated pricing game with $c_{1}=c_{2}=1$, evaluated at $t=50$ with $N_{\text{discrete}}=4,N_{\text{base}}=40$. [^11] We observe that all policies stop updating after about 30 rounds. Prices are discretized into four levels due to API costs, and initial states are uniformly sampled for the base games.
@@ 5 Discussion
Natural extensions of our framework include modeling heterogeneous beliefs on the opponent’s cost or representation initialization and analyzing through Bayesian Nash equilibrium, as well as broadening meta-strategy coverage, particularly for LLMs with richer prompt designs and pretraining histories. Additionally, cross-algorithm meta-games could shed further light on how algorithmic diversity affects the emergence and stability of collusion.
@@ A.2 Normalization of Q-values
$+\bigg(\tilde{Z}_{j}(s,p)-\min_{p^{\prime}}\tilde{Z}_{j}(s,p^{\prime})\bigg)\times\frac{V^{\pi_{j}|\pi^{\prime}_{j}}(s)-\min_{p^{\prime}}Q^{\pi_{j}|\pi^{\prime}_{j}}(s,p^{\prime})}{\max_{p^{\prime}}\tilde{Z}_{j}(s,p^{\prime})-\min_{p^{\prime}}\tilde{Z}_{j}(s,p^{\prime})}$ (10)
@@
After normalization, the original argmax of $\tilde{Z}_{j}$ is preserved: $Z_{j}(s,\pi_{j}(s))=V^{\pi_{j}|\pi^{\prime}_{j}}(s)+Q^{N}+f\cdot(Q^{M}-Q^{N})-\max_{s^{*}}V^{\pi_{j}|\pi^{\prime}_{j}}(s^{*})$, and the minimum satisfies $\min_{p^{\prime}}Z_{j}(s,p)=\min_{p^{\prime}}Q^{\pi_{j}|\pi^{\prime}_{j}}(s,p^{\prime})+Q^{N}+f\cdot(Q^{M}-Q^{N})-\max_{s^{*}}V^{\pi_{j}|\pi^{\prime}_{j}}(s^{*})$. The maximum Q-value (at the argmax action) maps to the true value function $V^{\pi_{j}|\pi^{\prime}_{j}}(s)$ and the minimum Q-value maps to $\min_{p^{\prime}}Q^{\pi_{j}|\pi^{\prime}_{j}}(s,p^{\prime})$ with an offset $Q^{N}+f\cdot(Q^{M}-Q^{N})-\max_{s^{*}}V^{\pi_{j}|\pi^{\prime}_{j}}(s^{*})$.
@@ A.3 Categorizing Q-learning Pretrained Policies via PC and CR
Figure 6: The PC (Def. 3.1) and CR (Def. 3.2) of 1,000 pretrained Q-learning policies with 4 and 15 discretized actions, respectively, under the experiment setting described in Sec. 4.2.
@@ A.6 Additional Results on Q-learning-based Meta-strategies
(b) Change in PC over time. Overall, PC improves over the course of test time. For RD strategies, although their own $V$ improves, their opponent’s does not, as the value against RD is already high at initialization. For both LC and RC, larger learning rates tend to benefit their opponents more, even when their own $\Delta(V^{\pi_{j}|\pi_{-j}})$ is comparable to that under smaller learning rates. This suggests that LC and RC with large learning rates provide opportunities for cooperation.
@@
Figure 10: Best-response graph for Q-learning with $c_{1}=c_{2}=0.8$, $f=1$ and $t=10,000$.
Figure 11: $\Delta_{t}$(CR) and $\Delta_{t}$(PC) for Q-learning with $c_{1}=c_{2}=0.8$, $f=1$ and $t=10,000$.
@@
Figure 12: Best-response graph for Q-learning under symmetric costs $c_{1}=c_{2}=1$ with optimistic initialization $f=1$ and a time horizon of $t=3,000$.
@@
Table 11: Average payoffs over 4,000 runs in the payoff matrix of Q-learning-based strategies under asymmetric costs, $c_{1}=1$ and $c_{2}=0.8$, and $t=10,000,f=1$. Standard errors are omitted as they are negligible in the order of $10^{-4}$. For reference, in this setting, the competitive and monopoly payoffs for player 1 are $\bar{r}_{c=1.0}^{N}=0.17$ and $\bar{r}_{c=1.0}^{M}=0.30$, and for player 2 are $\bar{r}_{c=1.0}^{N}=0.31$ and $\bar{r}_{c=0.8}^{M}=0.44$.
@@
Figure 14: Best-response graph for $c_{1}=1,c_{2}=0.8$, $f=1$ and $t=10,000$. The directed edge $(u\rightarrow v)$ corresponds to $c_{u}=1.0$ and $c_{v}=0.8$.
Figure 15: Best-response graph for $c_{1}=1,c_{2}=0.8$, $f=1$ and $t=10,000$. The directed edge $(u\rightarrow v)$ corresponds to $c_{u}=0.8$ and $c_{v}=1.0$.
Figure 16: $\Delta_{t}$(CR) and $\Delta_{t}$(PC). The high-cost strategies are plotted with fillings.
Symmetric costs $c_{1}=c_{2}=1$ with either optimistic or pessimistic initializations ($f\in\{1,0.5\}$) under time horizon $t=10,000$.
@@
Table 12: Average payoffs over 4,000 runs in the payoff matrix of Q-learning-based strategies under symmetric costs $c_{1}=c_{2}=1$ with either optimistic initialization $f=1$ or pessimistic initialization $f=0.5$. Time horizon $t=10,000$. For reference, in this setting, the competitive and monopoly payoffs are $\bar{r}^{N}=0.22$ and $\bar{r}^{M}=0.34$.
@@
Figure 17: Best-response graph for Q-learning with $c_{1}=c_{2}=1$, $f\in\{1,0.5\}$ and $t=10,000$.
Figure 18: $\Delta_{t}$(CR) and $\Delta_{t}$(PC) for Q-learning with $c_{1}=c_{2}=1$, $f\in\{1,0.5\}$ and $t=10,000$.